Grade 12th passPhysical Chemistry

Sir please tell me...what is the charge on 25 gram of carbonate ion....

Sheetal
8 Years agoGrade 12th pass

1 Answer

Sumit Raj pal
8 Years ago
25 g Carbonate ions .Molecular weight of CO3^-2 = 60 g/molno of mole = given wt/mol. Wt = 25/60 = 5/12 We know, in 1 mole of Carbonate ions , 2 moles of charge occur .So, in 5/12 mole of Carbonate ions , 10/12 mole of charge occur .So, charge = 10/12 × 6.023 × 10²³ × 1.6 × 10^-19 C = 8.03 × 10⁴ C